(1-x^2)d^2y/dx^2 -2x dy/dx+n(n+1)y=0


1
Expert's answer
2021-07-26T14:37:51-0400

"=" "(1-x^{2})\dfrac{d^{2}y}{dx^{2}}-2x\dfrac{dy}{dx}+n(n+1)y=0" .........................................1)


This equation is called LEGENDRE'S polynomial equation .


substituting -

"=y=a_0x^m+a_1x^{m+1}+a_2x^{m+2}+.....(a_0\neq0),"


Now , series first take of the form -


"=a_0m(m-1)x^{m-2}+a_1(m+1)mx^{m-1}+.......[a_{r+2}(m+r+2)(m+r+1)-{(m+r)(m+r+1)-n(n+1)}a_r]x^{m+r}+....=0"


Equating to zero the co-efficient of the lowest power of "x" , i.e. of "x^{m-2}," we get -


"=a_0m(m-1)=0, m=0,1" "[\because" "a_0\neq0]"


Equating to zero the coefficient of "x^{m-1}\ and \ x^{m-r}\ , we" get "a_1(m+1)m=0.....................2)"



"a_{r+2}(m+r+2)(m+r+1)-{(m+r)(m+r+1)-n(n+1)}a_r=0...............................(3)"



When m=0, (2) is satisfied and therefore ,"a_1\neq0." Then (3) gives , taking r=0,1,2........in turn ,


"a_2=\dfrac{-n(n+1)}{2}a_0" ", \ a_3=\dfrac{(n-1)(n+2)}{3!}a_1"


"a_4=\dfrac{-(n-2)(n+3)}{4.3}a_2=\dfrac{n(n-2)(n+1)(n+3)}{4!}a_0"


"a_5=\dfrac{-(n-3)(n+4)}{5.4}a_3=\dfrac{(n-1)(n-3)(n+2)(n+4)}{5!}a_1, etc"


Hence for m=0 , there are two independent solutions of (1),


"y_1=a_0({1-\dfrac{n(n+1)}{2!}}x^{2}+\dfrac{(n-2)n(n+1)(n+3)}{4!}x^{4}-............4)"



"y_2=a_1(x-\dfrac{(n-1)(n+2)}{3!}x^{3}+\dfrac{(n-3)(n-1)(n+2)(n+4)}{5!}x^{5}-.....)" ......................5)


When m=1 , (2) shows that "a_1=0." Therefore (3) gives ,


"a_3=a_5=a_7=0"


"and" "a_2=\dfrac{-(n-1)(n+2)}{3!}a_0"



"a_4=\dfrac{(n-3)(n-1)(n+2)(n+4)}{5!}a_0, etc"


Thus for m=1 , we get the solution of (5) again , Hence "y=y_1+y_2" is a general solution of (1).


So our general solution of given differential equation is "y_1+y_2"